OpenStudy (anonymous):

suppose that a manufacturer's weekly profit in dollars is given by P= -2x raise to the second power + 240x , how many items x must be produced to realize a profit of 5,4oo pesos ?

OpenStudy (anonymous):

\[-2x^2+240x=5400\] \[-2x^2+240x-5400=0\] \[x^2-120x+2700=0\] \[(x-90)(x-30)=0\] \[x=90\text{ or } x=30\]

OpenStudy (anonymous):

is this the right answer ?

ganeshie8 (ganeshie8):

the answer is as right as the question :)